Leg 210 began at 1712 hr on Sunday, 6 July 2003, when the ship arrived in St. Georges, Bermuda. As the ship arrived 1 day early, the planned berth at the Royal Naval Dockyard on the west end of the island was not yet available. Immigration and custom clearance was completed that day. Because of the change in berth location, only one ODP container and the airfreight could be loaded on Monday, 7 July. The drill crew continued to work on replacing the drill line; unfortunately the new drill line reel was larger, necessitating significant reworking of the mounting structure.
One Tuesday, 8 July the ship moved to the Royal Naval Dockyard (RND) on the west end of Bermuda. After the pilot boarded (0745 hr), all lines were released (0824 hr), and the tug pulled the ship clear of the harbor (0831 hr), the ship began the transit to the RND. The first line was ashore at the RND at 1035 hr. The early arrival in Bermuda and berth change resulted in port call activities not effectively starting until 8 July.
The ODP crew change was completed at 1430 hr on 8 July. This crew change was extended so that the technical staff could be informed about the status of IODP and changes in medical and retirement benefits. The Lawrence Berkley National Labs X-ray computed tomography imaging (CAT) scanner was installed, and testing, calibration, and training began. Three containers of fresh groceries were loaded on Tuesday. Although the reefers had been left without power by dockworkers for ~24 hr, no spoilage occurred. ODP offloaded three 20-ft containers of supplies. We began loading 1400 metric tons of fuel on 8 July. Due to the small transfer hose size, this took 2 days; it shut down all drill floor activities for the drill line replacement because of welding restrictions. Potable water was delivered to the ship by small truck on an hourly basis.
Other activities on 8 July included the following:
Activities related to troubleshooting AHC problems continued on Wednesday, 9 July with a technical review meeting with the Leg 209 ODP electronics technician, the Transocean operations manager, and the vendor service engineers. The rental MRU was mounted and the existing MRU was diagnosed with severe calibration problems.
On 9 July the ship had to shift down the dock to make way for another arriving ship. The pilot was on board at 0915 hr, and the lines were cleared at 0950 hr; a tug was tied to the stern, and we started shifting along the pier. At 1020 hr, the vessel was secured to the wharf two ship lengths down the pier. Testing continued on the DSS and the RMM as well as the WOB filter. The departing ODP dry ice and foreign airfreight shipments were sent out and all departing ODP surface freight was staged on the dock. One of the drill collar (DC) racks was cleared out and then loaded with eight DCs and three tapered drill collars (TDCs). Delivery to the ship of the 30-ft DCs was difficult because of the very narrow roads in Bermuda. Two flats of DCs were placed on the dock in preparation for loading. American Bureau of Shipping surveyors were on board carrying out various required ship inspections. More tours were conducted for staff of the Bermuda Biological Station and Texas A&M University Research Foundation representatives.
The Transocean crew crossover occurred on Thursday, 10 July. The new drill crew reorganized the DC rack and loaded the remaining 16 DCs. The reentry cone was spotted in the moonpool for assembly. The drill line installation was completed at 2100 hr. The AHC vendor engineers began testing and calibrating the AHC and then provided training for the drillers regarding set up and operating procedures for the AHC. The remaining ODP airfreight arrived on the dock on late Thursday evening, and the departing ODP containers were sealed for shipment.
On Friday, 11 July, the Leg 210 airfreight was loaded. The pilot and tug were ordered for a 1500 hr departure. A comprehensive safety equipment and lifeboat orientation was held for new oncoming personnel and the science party.
The pilot boarded at 1631 hr on Friday 11 July, the tug boats secured to the ship at 1638 hr, and the last line was released at 1648 hr. Once the tugs were clear of the ship (1655 hr) we transited to the pilot station with beautiful blue skies and calm seas. The pilot departed the ship and we began the transit to Site 1276 (proposed Site NNB-01A) at 1813 hr. The Schlumberger logging engineer boarded the ship via the pilot vessel on 11 July.
During the transit from Bermuda, we assembled the reentry cone, picked up and determined the space-out of the bottom-hole assembly (BHA) for drilling in the 20-in casing, and performed general rig maintenance. On 14 July, we crossed the Gulf Stream and entered dense fog and cooler weather. That evening, we slowed to 9.5 nmi/hr as we entered marked fishing grounds in fog. During the final approach to Site 1276, we had to change heading slightly to avoid some research vessels that were towing gear. After a transit of 1227 nmi at an average speed of 11 nmi/hr, we arrived at Site 1276 at 0800 hr on Wednesday, 16 July.
Once we arrived at Global Positioning System (GPS) coordinates of Site 1276 (45°24.23'N, 44°47.15'W), the thrusters were lowered and switched to dynamic positioning (DP) mode at 0815 hr on 16 July. A seafloor positioning beacon was deployed at 0844 hr. The precision depth recorder indicated a water depth of 4564.5 meters below sea level (mbsl) (4575.4 meters below rig floor [mbrf]).
Figure F32 shows a diagram of drilling, casing, and coring in Hole 1276A, and Figure F33 gives a breakdown of operations activities while on site.
During the transit we had prepared the 20-in casing running tools (CRTs), BHA components, pilot bit, underreamer, centralizer, casing shoe, and reentry cone. The underreamer arms were set for 22 in. After the ship was on location in DP mode, we assembled DCs into five stands and stored them in the derrick. The upper guide horn was removed and stored in the DC rack so the reentry cone could be centered over the moonpool.
A DC was attached to the top of the CRT, and the CRT was attached to the 20-in casing hanger and secured in the derrick. At 1430 hr on 16 July, we started rigging up for assembling eight joints of 20-in casing (total length = 101.9 m below the reentry cone mud skirt). The casing shoe joint was positioned in the rotary table and the casing assembled; the lower three casing connections were spot welded. Once the casing hanger was attached to the top, we lowered the 20-in casing through the reentry cone and latched the hanger into it. We lifted the entire casing and reentry cone assembly to ensure it was securely attached and to determine its weight (casing = 24,000 lb; casing + reentry cone = 38,000 lb). We then disconnected the CRT from the casing hanger and welded the casing to the reentry cone. The CRT, with two joints of DCs attached to the top, was stored in the derrick.
Our next step was to assemble the drill-in BHA. At 1930 hr, the drill crew started assembling a BHA consisting of an 18-in bit, bit sub with float, underreamer, crossover sub, mud motor, 14-in stabilizer, and 10 DCs to allow the bit to extend 4.83 m below the bottom of the casing. Once the mud motor was attached above the bit and underreamer, we flow tested this setup to ensure the mud motor and underreamer were functioning properly.
The drill-in BHA (total length = 106.73 m below reentry cone mud skirt) was lowered through the reentry cone and 20-in casing, and the CRT was attached to the top and latched into reentry cone at 0105 hr on 17 July. After visually inspecting the latching of the casing to the reentry cone and the CRT to the casing hanger, the entire assembly was raised to confirm that it was securely attached and to obtain a total weight (casing = 25,000 lb; casing + reentry cone = 40,000 lb; drill-in BHA below the CRT = 40,000 lb; BHA above CRT = 10,000 lb). We lowered the reentry assembly through the moonpool (0205 hr) and prepared the rig floor for running the assembly to the seafloor. As the assembly was lowered to 2701 mbrf, we filled the drill pipe with seawater every 10 stands.
Because of the deep water at Site 1276 and planned penetration depth, we had to use some drill pipe that had not been used for a year. This resulted in some extra time expended to remove pipe thread protectors that were difficult to remove and to clean the threads. Because of rust inside the drill string, each time we picked up a stand we had to run a rabbit through the pipe; we used a rabbit with a smaller outside diameter than normal to ensure the inside of this drill string was open.
To minimize the risk that a large quantity of rust particles could be flushed down the drill string, potentially damaging the mud motor, we circulated a complete drill string volume of seawater through the drill pipe when the bit was at 2701 and 3621 mbrf. We deployed the VIT camera system and continued to lower the drill string. We then picked up the top drive, spaced out the drill string, and established circulation; mud motor circulation rates were recorded just off bottom.
Based on a reduction in drill string weight, the driller tagged seafloor at 2045 hr on 17 July, initiating Hole 1276A. The seafloor depth as felt by the drill bit was 4563 mbrf (4552.1 mbsl), but this was later corrected to 4560 mbrf (4549.1 mbsl) (Table T2), as the reentry cone landed and could not be lowered past this depth. It took a total of 28 hr to drill the 20-in casing to 4662.0 mbrf (102.0 mbsf).
We experienced highly variable penetration rates due to the presence of several hard layers and, at times, what appeared to be a plugged jet in the bit. Drilling times were mostly between 2 and 40 min/m, except for some very hard layers at 16, 56, and 62 mbsf that took 150, 80, and 210 min/m, respectively. We infer that these were intervals containing glacial dropstones. In general, the driller was using 100140 spm (500700 gpm) while drilling, with a resulting torque of 5000 ft·lb at the bit with a bit rotation of ~70 rpm. Penetration slowed again while drilling the last few meters, and the reentry cone appeared to land on the seafloor 3 m shallower than expected.
After drilling in the casing (completed at 0045 hr on 19 July), our next step was to activate the CRT to release the drill string from the 20-in casing and reentry cone. The driller reduced the string weight by 40,000 lb (the neutral point for the casing at the CRT) and then attempted to release the CRT at 0230 hr. Although it should have released relatively quickly, the CRT finally released at 0500 hr, when the driller rapidly increased the WOB. Even then, it still required repeated raising and lowering of the drill string to fully release the BHA from the casing. This sticking may have been due to the underreamer arms not being completely retracted or because the CRT was sticking in the casing hanger. Using the subsea camera system, we observed that sediment was completely covering the reentry cone. To better visualize the cone and seafloor, we retrieved the camera to change the guide funnel so it could pass over the CRT and reach down to the bit. During camera retrieval, we circulated seawater to attempt to wash away some of the sediment in the core. When the camera was redeployed and had passed over the CRT, we reconfirmed that the reentry cone was covered by sediment, although a "crater" in the sediment could be clearly imaged on the sonar. It appeared that we may have set the reentry cone base slightly below the seafloor. We raised the bit until it was just inside the reentry cone and jetted with seawater to try to clear the sediments. The bit cleared the seafloor at 1000 hr.
Next, we retrieved the camera system and then pulled the drill string out of the hole. The drilling assembly was then disassembled (DC stands stored in the derrick, CRT detorqued, and mud motor and underreamer flow tested and then flushed with fresh water). The bit cleared the rotary table at 2300 hr.
Once the underreamer and bit were back on the rig floor, we observed that the underreamer's nozzles were damaged. The damage was likely due to plugging of the nozzles as well as to clay and glacial dropstones packed inside the bit and underreamer. The sediment (late Pliocene age, uppermost part of fan deposits) and rocks were passed to the scientific party. The underreamer was subsequently repaired for drilling in the 16-in casing.
Our next operation was to install 16-in casing to 750 mbsf. The first step was to drill a 21-in hole to 800 mbsf. We started assembling the BHA at 0115 hr on 20 July. The BHA consisted of a 97/8-in wobble (pilot) bit, bit stabilizer, a 21-in bicenter reamer, a float sub, eight DCs, the DSS (which measures downhole WOB, torque, and pressure), 10 DCs, and a TDC. The 174-m-long BHA weighed 50,000 lb, allowing a maximum WOB of 35,000 lb. The bicenter reamer could pass through a diameter of 18 in and was capable of drilling a 21-in hole beneath the 20-in casing.
While lowering the BHA to the seafloor, the subsea camera system was deployed. At 1245 hr (with the bit at 4539 mbsl), the drill string length was adjusted in preparation to reenter Hole 1276A. The bit reentered the hole at 1300 hr. The reentry cone was buried in the sediment and only visible as a dark conical depression in the seafloor; the sonar could clearly image the sedimentary crater above the cone.
The top drive was picked up and we tagged the bottom of the at hole 4666.9 mbrf (106.9 mbsf). We alternated between passive and active heave compensation while drilling to check recent AHC calibrations and drilling procedures (detailed engineering report is available from ODP).
The rate of penetration (ROP) varied from 21 to 3 m/hr, generally decreasing downhole. Substantial reductions in ROP occurred at 503 mbsf (from 13 to 7 m/hr) and at 685 mbsf (from 7 to 34 m/hr). A failed connection in the top drive circulation system interrupted drilling for 3 hr when the bit was at 666 mbsf (0915 hr on 23 July). The bit was raised off the bottom (to 5165 mbrf [605 mbsf]) to repair the top drive. After the repair, we drilled back to bottom (609666 mbsf), encountering 1015 m of soft fill.
The wobble bit with bicenter reamer drilled smoothly with no indication of drag along the borehole walls. We circulated sepiolite mud (20 bbl) every 50 m while drilling. The bit reached 800 mbsf at 2330 hr on 24 July.
We circulated 50 bbl of sepiolite mud to clean the hole and then raised the bit back up to the bottom of the 20-in casing shoe (102 mbsf). No tight sections of hole were encountered on the way up. While lowering the bit back down, the bit encountered some resistance from 615 to 634 mbsf. We washed from there down to 800 mbsf and had to wash 15 m of soft fill from the bottom of the hole. We circulated another 40 bbl of mud to clean out the hole and then put 375 bbl of mud in the bottom of the hole so that mud would cover the tight interval at 615634 mbsf. As the drill string was raised, additional tight spots were found at 576432, 423412, and 288103 mbsf and they required use of the top drive. Once the bit was at the reentry cone and before pulling completely out, we circulated seawater to clean sediments from the reentry cone in order to make it easier for us to find the cone on the next reentry. The bit cleared the seafloor at 2100 hr on 25 July.
The BHA was taken apart and the bit was on the rig floor at 0640 hr on 26 July. The DSS was moved near the downhole laboratory so the data could be downloaded. The wobble bit and bicenter reamer had experienced 84 hr of rotation with 25,00035,000 lb WOB (7075 rpm) and could not be reused because of loose bearings.
Another problem with the top drive's seawater circulation system was observed when tripping out of the hole; water was leaking between the swivel and top drive shaft while circulating. At 0645 hr on 26 July, we started troubleshooting the top drive and swivel. The top drive was repaired, pressure tested to 1000 psi, and back in service by 1330 hr. Finally, preventative maintenance had to be done on the drill line (slip and cut).
We started preparing the rig floor for assembling the 16-in casing at 1515 hr. Once we were finished assembling the casing (749.11 m; 0200 hr on 27 July) the CRT was attached to the casing hanger and the casing was lowered below the rig floor and hung off on the moonpool doors. The CRT was disconnected, and we started assembling the 16-in drilling BHA through the 16-in casing. The BHA consisted of a 14-in tricone bit, underreamer (arms set to 20 in), a mud motor, a 14-in centralizer/stabilizer, 12 DCs, and 21 stands plus a double of drill pipe. The BHA length was 753.90 m, so the pilot bit extended 4.83 m below the 16-in casing shoe.
We attached the CRT to the top of the BHA (1115 hr), lowered it to the moonpool, and attached the CRT to the 16-in casing hanger. We picked up the entire drilling/casing assembly (1245 hr) and deployed the camera system to inspect the casing as well as to confirm the spacing of the bit and underreamer below the casing.
At 1415 hr we began lowering the assembly to the seafloor. The bit was at the seafloor and ready for reentry at 0000 hr on 28 July. After only 10 min (0010 hr), we reentered Hole 1276A. When the bit was at 4649 mbrf (89 mbsf; 0100 hr), we attached the top drive. The bit passed through the base of the 20-in casing (102.07 mbsf) at 0200 hr.
From the 20-in casing shoe to 4930.79 mbrf (370.79 mbsf), the drilling time for each 9.6-m interval was 312 min. The drilling parameters for this section were WOB = 414 klb, circulation rate = 80110 spm, and pump pressure = 6501200 psi. We started circulating 30 bbl of mud every stand starting at 284 mbsf. We continued drilling the casing in from 5017 to 5050 mbrf (457490 mbsf; WOB = 2030 klb, circulation = 150 spm, pressure = 2100 psi) until 0715 hr on July 29; at this time, it became apparent that the mud motor was not working properly. As we pumped 30 bbl of mud the pump pressure increased by 700 psi. Although this was not typical for a mud motor stall, we could not restart the mud motor despite repeated attempts at lifting off bottom or varying the circulating rate and WOB. At 1000 hr, we decided to retrieve the drill string to replace the mud motor.
The bit cleared the seafloor at 1415 hr, and the entire drilling/casing assembly was raised back to the ship. We hung the 16-in casing on the moonpool doors (2300 hr) and then began to take apart the BHA (2330 hr). Before taking apart the mud motor, underreamer, and bit, we tested the equipment by pumping seawater through it. The mud motor had seized up and would not rotate. Furthermore, the underreamer arms would not open and it was packed off with sand (including sand under the piston that actuates the arms). Later, we found pieces of rubber jammed in the nozzles of the underreamer; these were from inside the mud motor, further documenting the failure of its internal mechanism.
The underreamer had to be completely taken apart and nearly all moving parts were replaced. At 0430 hr on 30 July, we assembled the pilot bit, the rebuilt underreamer, and a new mud motor. Once this was tested, we started to reassemble the drilling BHA (0900 hr). The BHA was latched into the 16-in casing hanger at 1130 hr, and we started lowering the drilling BHA and casing back down to the seafloor.
Once the bit was near the seafloor (2000 hr), it took 45 min to reenter Hole 1276A. The bit was lowered into the hole to 92 mbsf when we picked up the top drive to prepare for drilling. We began drilling the 16-in casing into the hole at 2200 hr. Drilling in the casing went quite quickly and smoothly down to 389.87 mbsf (drilling time = 214 min/9.6 m, WOB = 212 klb, circulation = 145155 spm; pressure = 19002200 psi). We encountered what appeared to be a series of ledges at 392, 394, and 396 mbsf. It took 120 min to drill the interval from 389.87 to 399.47 mbsf. After each ledge was drilled we had to use 20 klb to work the casing past the ledge. While passing through this interval, the mud motor appeared to be experiencing torque, indicating that the underreamer arms were cutting hole. It is also possible that the underreamer arms were exposed to shock loads (perhaps as high as the full casing/BHA weight) when the casing would finally break through a ledge. From 389.87 to 754.05 mbsf, drilling time was 25120 min/9.6 m; circulation was 155160 spm; and pump pressure was 21502300 psi. Several hard intervals were encountered at 698, 703, and 743 mbsf. The casing hanger landed in the reentry cone at 1115 hr on 2 August with the bottom of the 16-in casing at 749.07 mbsf.
The 16-in casing landed 2.8 m higher than expected. We interpret that this was due to stretching of the drill pipe under the very heavy weight of the casing and drilling BHA; this is consistent with drill pipe stretch calculations. When we had deploying the 20-in casing and reentry cone, we calculated a drill pipe stretch of ~9.6 m. With the longer, much heavier 16-in casing string and drilling BHA, the calculated stretch was 12.4 m.
After landing the casing, it took only 30 min to release the CRT (1145 hr). We then reduced the circulation rate, pulled the bit and underreamer back into the casing, and started pulling out of the hole. The bit cleared the seafloor at 1440 hr. Before we could retrieve the drill string, we conducted routine preventative maintenance on the drill line (slip and cut). After the mud motor, underreamer, and pilot bit were flushed out with fresh water, the bit arrived back on the rig floor at 0340 hr on 3 August.
We observed that the underreamer was missing its three outer cones (cutting structure on the arms) as well as two sets of pins that control how far the arms extend. These were left in the hole. The body of the underreamer and bit exhibited some deep gouges, suggesting that the cones had been caught between these and the formation. We also observed water leaking from the hard covering over the top of the underreamer's bottom connection. Even after losing their cones, the underreamer arms appear to have continued cutting the formation because the outer parts of the arms had been worn back.
After opening the hole to 800 mbsf and drilling the 16-in casing in to 749.07 mbsf, our next operation was to cement the casing in place. We had originally planned to use a cement retainer. Because this would have required an extra pipe trip, we decided to use a 16-in Cameron cup tester (CCT) for the cementing operation instead.
At 0345 hr on 3 August we tested the rig floor cementing equipment and then assembled the cementing BHA. The BHA consisted of a short piece of drill pipe, the CCT, an 8-in DC, a three-blade stabilizer, four DCs, a TDC, and five joints of 51/2-in drill pipe (length = 110.09 m; weight = 27,000 lb).
We lowered the BHA and were ready for reentry at 1400 hr. The ship was positioned precisely at the same location where we had efficiently reentered before and where each time the bit was directly over the hole. However, this time there was no indication of the location of the hole on either the camera or sonar.
After spending 1 hr searching for the hole with no success and with the ship directly over the previous reentry coordinates, we decided to probe the seafloor with the drill string in an attempt to locate the hole. The camera imaged a circular shadow, and we lowered into it at 1455 hr. It appeared that we had entered the hole because there was no reduction in drill string weight, and the drill string was lowered to 4643 mbrf (83 mbsf). At this point, we observed on the camera image that the drill string did not appear to be moving down into the seafloor. We immediately stopped lowering and raised it back up to 4569 mbrf (9.4 mbsf). We slowly lowered the drill string again and it started meeting resistance at 13.4 mbsf. At this point, we decided to circulate seawater to attempt to clear sediment from the area. We also decided to raise the camera and inspect the drill string for damage. Unfortunately, we observed that the TDC was severely bent, so we had to retrieve the drill string before continuing. First, we conducted a sonar search of the area. We saw a sonar return ~60 m from the bit location, but we did not move closer to investigate. At 1900 hr, we began retrieving the drill string. When the end of the drill string was at 107 mbrf (0230 hr on 4 August) we had to cut off the connection from the last joint of 51/2-in drill pipe and we removed the bent (~20°) TDC.
We replaced the TDC and started lowering the BHA back to the seafloor at 0400 hr. This time we attached a transponder to the camera/sonar system frame so that its position (and the end of the drill string) could be accurately monitored.
When the drill pipe was ready for reentry (1230 hr; 4554 mbrf, or 6 m above seafloor), we observed that the end of the drill string was offset 5570 m from the moonpool. This was caused by a significant increase in bottom-current strength since the previous reentries. Moving the drill pipe to the location where it had just previously been lowered into the seafloor, we observed in the camera image a ~1-m-diameter hole in the seafloor together with an adjacent smaller hole. After discussion, we decided that this marked the previous attempted reentry and it was not the location of the Hole 1276A cone. At this time, we began seeing plumes of sediment-laden water flowing from the southwest. We decided to trace these to their source, where we found a depression in the seafloor that turned out to be the sedimentary crater in Hole 1276A. The rim of the depression was visible on the sonar. Sediment-laden water was streaming from the seafloor around this depression, which was ~15 m in diameter. The next challenge was to lower the drill string into the 3.5-m-diameter cone that was obscured by sediment and flowing muddy water and had a 75-m offset between the moonpool and the bit position.
During our first reentry attempt (1920 hr), the CCT appeared to hang up on the throat of the reentry cone (4563 mbrf; 3 mbsf). We slowly lowered the drill string to 4589 mbrf (29 mbsf) where it encountered an obstruction. We started circulating seawater to wash down from 4589 mbrf (29 mbsf), but tagged something hard at 4594 mbrf (34 mbsf; 0200 hr on 5 August). It was clear that the CCT was hanging up in the reentry cone, so we pulled out of the seafloor at 0325 hr.
After ~1 hr of repositioning the drill pipe directly over the hole, we attempted another reentry (0522 hr). Once again, we could not lower past 4564.5 mbrf (4.5 mbsf) in the cone. After 2 hr of trying we decided to pull the pipe clear of the reentry cone (0730 hr) and cease attempts to use the CCT for cementing. The CCT reached the rig floor at 1630 hr. A total of 61 hr was expended while attempting reentry for the cementing job.
We decided to abandon attempts to cement the casing, as we felt that by now the formation had likely collapsed around the outside of the 16-in casing and additional time to cement would create even more risk to achieving any of the deep, primary leg objectives. We felt that it was important to verify that reentering Hole 1276A was possible and that the smaller 97/8-in RCB bit was the best tool to use. Furthermore, we still did not know what impact the pieces of the underreamer previously lost in the hole would have on our ability to advance the hole.
After preparing the rig floor (1830 hr on 5 August), we started assembling the BHA. The BHA consisted of a rotary core barrel (RCB) bit (C3), a mechanical bit release (MBR) (so we could log after coring), one DC, the DSS (to measure downhole WOB, torque, and pressure), 13 DCs, a TDC, and 6 joints of 51/2-in drill pipe (length = 202.08 m; weight = 60,000 lb). When the bit was at 4550 mbrf (0500 hr on 6 August), we had to do drill line maintenance (slip and cut) before we started the coring operations. At 0630 hr, we started reentry attempts in Hole 1276A, where sediment-laden fluid was still billowing out of the seafloor around the hole. After almost 5 hr, we succeeded reentering the hole on our fourth attempt (1125 hr).
We lowered the drill pipe to 4.7 m below the 16-in casing (5313 mbrf; 754 mbsf), where the bit encountered some resistance. A core barrel (wash core) was deployed and we drilled down to 5360 mbrf (800 mbsf; WOB = 10 klb, bit rotation = 70 rpm; circulation = 85 spm at 1100 psi, torque = 150 A). After recovering the wash core (Core 1W; 2355 hr), we began RCB coring at 5360 mbrf (800 mbsf). By 0025 hr on 10 August, we had cored from 800 to 1059.7 mbsf and recovered 210 m of core (recovery = 81%). The time to cut each core varied primarily from 22 to 95 min (average = 50 min) except for Core 9R (876.8886.4 mbsf) which took 120 min. Hole conditions remained good throughout the cored interval.
Coring parameters started at WOB = 12 klb, bit rotation = 70 rpm, and circulation = 85 spm at 1100 psi, but by Core 28R were WOB = 2025 klb, bit rotation = 6070 rpm, and circulation = 115 spm at 1800 psi. Thirty barrels of sepiolite mud was circulated every core. Particulate tracers for microbiological experiments were deployed while cutting Cores 8R, 18R, 19R, and 28R. The RMM was run on top of the core barrel while taking Cores 4R, 7R, 13R, and 24R to download data from the DSS. The AHC was used while cutting Cores 14R, 15R, 16R, 18R, 20R, and 22R. While cutting Core 22R, it developed a hydraulic leak, which was subsequently repaired.
RCB coring continued from 1059.7 to 1338.1 mbsf. Cores 29R to 57R penetrated 278.4 m of section and recovered 245.75 m (recovery = 88%). Coring parameters for Cores 29R through 32R were WOB = 20 klb, bit rotation = 70 rpm, circulation = 110 spm at 1600 psi, and torque = 175200 A. During a wiper trip after Core 32R, we encountered no drag from the base of the hole (1098.2 mbsf) up to the 16-in casing at 749 mbsf. On the way back down to the bottom of the hole, tight spots had to be reamed out at 803, 864, and 874 mbsf. After 10 m of soft fill on bottom had been cleaned out, we pumped 50 bbl of sepiolite mud to clean cuttings out of the hole. After the 6.5-hr wiper trip, we resumed coring. From 1615 hr on 10 August to 0605 hr on 13 August, we cut Cores 33R to 57R (1098.21338.1 mbsf) and recovered 212.15 m (recovery = 88%). ROP varied from 19 to 6.4 m/hr and generally decreased downhole. Sepiolite mud (30 bbl) was circulated after each core. Coring parameters were WOB = 20 klb, bit rotation = 70 rpm, torque = 175 A, and circulation = 110 spm at 16001700 psi until the lower part of this section where WOB, torque, and pump rates started to increase (WOB = 25 klb, torque = 200300 A, and circulation = 130 spm at 2300 psi, respectively). Particulate tracers for microbiological studies were deployed on Core 48R.
There was a significant decrease in recovery in Cores 55R to 57R. Recovery had been averaging 93% in Cores 33R to 54R, but it dropped to 30%60% in Cores 55R to 57R. Torque and pump pressure also increased while cutting Core 57R (torque = 400 A and pressure = 2600 psi, respectively). When Core 57R was recovered, we immediately noticed that its diameter was significantly reduced. Although the bit had only 44.8 cutting hours, the sudden change in core quality indicated that one or more of the core guides may have been bent inward or perhaps a single cone was not rotating properly. We decided to stop coring because we wanted to minimize the chances that a bit failure might junk the hole, and we prepared to log the hole.
Before we conducted a wiper trip to prepare for logging, we circulated 60 bbl of sepiolite mud to clean cuttings out of the hole and we then raised the bit up into the 16-in casing (07301030 hr on 13 August). We lowered the bit into the hole to 5626 mbrf (1066.8 mbsf; 10301215 hr), where we used the top drive to drill some tight spots down to 1095.6 mbsf. Once this section was clean, we removed the top drive and we were able to freely lower the bit to 1297 mbsf. At 1530 hr, we had to use the top drive again to drill from 1297 to 1338 mbsf. The bit was back at the bottom of the hole at 1630 hr. We circulated 50 bbl of mud and then displaced the hole with 185 bbl of mud before starting to pull out of the hole to drop the bit.
We raised the bit up with the top drive in place up to 1297 mbsf, where it was racked back. While we continued to pull out of the hole, the subsea camera/sonar system was deployed at 2015 hr so we could observe bit release and reentry for logging. The bit cleared the seafloor at 2210 hr, the ship was offset 50 m to the south, and at 2245 hr we lowered the rotary shifting tool (RST) on the wireline to release the bit. The bit did not release and we lowered the RST again to close the MBR ports. When the sleeve was shifted back in place over the ports, at 0020 hr on August 14, we observed the bit falling off onto the seafloor. At 0130 hr, the drill pipe was adjusted for reentry at 4553 mbrf (7 m above the seafloor).
At 0130 hr the DP operators began positioning the vessel for reentry. The offset between the beacon on the camera/sonar frame and the moonpool was 2540 m. When the end of the pipe passed over the hole, we observed sediment-laden fluid still coming out of the hole and obscuring it. The sedimentary crater was ~15 m in diameter. After an initial failed stab at the hole, we successfully reentered Hole 1276A at 0400 hr.
The end of the pipe (EOP) was lowered to 186.6 mbsf, and we retrieved the camera/sonar system. At 0445 hr, we started preparing the rig floor to run the wireline logging tools. The camera system was back on deck at 0630 hr and we started assembling the triple combination (triple combo) tool string. Unfortunately, once the logging tools were deployed, they would not pass an obstruction ~5 m below the base of the 16-in casing (749 mbsf). We retrieved the logging tool and decided to try lowering the open EOP past the obstruction. We lowered the EOP to 821 mbsf (~21 m into the 97/8-in RCB hole) and deployed the triple combo a second time. Once again, the tool did not pass more than 3 m out the end of the pipe. It was now clear that the logging tools could not get into the open hole, so we abandoned attempts to log and started to pull out of the hole to resume RCB coring. The triple combo tool was back on the rig floor at 2020 hr, and we began retrieving the drill string at 2230 hr. The EOP cleared the seafloor at 0020 hr on 15 August and was back on the rig floor at 0840 hr.
After attaching a new RCB bit, a new MBR, and removing a damaged piece of 51/2-in drill pipe, we lowered the coring BHA down to the seafloor. The subsea camera/sonar system was deployed at 1730 hr when the bit was at 4517 mbrf (42 m above seafloor). Before we could reenter, we conducted some preventative maintenance on the drill line (slip and cut); this took from 1800 to 1930 hr.
At 1930 hr, we began positioning for reentry with the bit at 4553 mbrf. We could clearly see the rim of the sediment crater above Hole 1276A on the camera image, but this time the hole was not shrouded by sediment-laden water. We reentered Hole 1276A on the second attempt at 2145 hr and lowered the bit to 736 mbsf. After retrieving the subsea camera/sonar system, we started lowering the bit into the open hole below the 16-in casing (0015 hr on 16 August). The bit encountered as much as 25 klb of drag, so we installed the top drive to drill and circulate back to the bottom of the hole at 1338.1 mbsf. Tight spots in the hole had to be worked at 825, 834839, 877882, 904, 1005, 1053, 1085, 1148, 1170, 1236, and 1254 mbsf. These zones required WOB = up to 10 klb, bit rotation = 75 rpm, torque = 175300 A, and circulation = 100 spm at 1100 psi. Thirty barrels of mud were circulated at 1120 mbsf. Fill had to be drilled and washed from the hole at 1264.61338.1 mbsf. Before coring, we circulated 50 bbl of mud at 1338 mbsf.
We started coring at 1330 hr on August 16. Cores 58R to 61R were taken from 1338.1 to 1372.9 mbsf, and we recovered 32.66 m (recovery = 93%). ROP was 10.2 m/hr, and coring parameters were WOB = 1520 klb, bit rotation = 70 rpm, circulation = 120 spm at 2300 psi, and torque = 200300 A. After cutting each core, 30 bbl of mud was circulated.
Cores 62R to 80R were cut from 1372.9 to 1555.4 mbsf (0225 hr on 17 August to 1450 hr on 19 August) and we recovered 170.27 m (recovery = 93%). Drilling parameters remained relatively constant in this interval: WOB = 2025 klb, bit rotation = 70 rpm, torque = 200250 A, and circulation = 120130 spm at 23002500 psi. The ROP decreased downhole from 9.3 m/hr (Cores 62R to 70R) to 6.2 m/hr (Cores 71R to 77R) to 3.8 m/hr (Cores 78R to 80R). The AHC was used while cutting Core 64R and then used for all odd-numbered cores starting with Core 67R.
A full-length nonmagnetic RCB core barrel was deployed for all odd-numbered cores, and we continued to use a standard barrel for even-numbered cores. Particulate tracers for microbiological contamination testing were deployed every tenth core starting with Core 69R.
The bit encountered a hard interval at 1412.4 mbsf. A substantial reduction in penetration rate occurred between Cores 75R (85 min to cut 9.6 m; ROP = 6.7 m/hr) and Core 76R (170 min to cut 9.6 m; ROP = ~3.3 m/hr). This slow rate of penetration persisted while cutting Cores 77R to 80R. These cores also had slightly lower recovery. The recovered cores showed no obvious reason for the substantial change in penetration rate.
Because of the slow ROP and consistent torque, we decided to conduct a wiper trip up to the 16-in casing. At 1600 hr on 19 August, we circulated 50 bbl of sepiolite mud and then started raising the bit. The top drive was used to rotate and circulate as the bit was raised to 1284 mbsf. At 1830 hr, the top drive was removed and the bit was raised the rest of the way up to the 16-in casing. The wiper trip to this point lasted 5.75 hr.
We began lowering the bit back into the hole at 2045 hr. The bit immediately ran into a tight area at 851.5 mbsf, and we had to install the top drive to be able drill through it. Drilling did not start until 0000 hr on 20 August because we first had to replace a damaged 20-ft piece of drill pipe ("knobby") at the rig floor.
It took 13.5 hr to drill back to the bottom of the hole at 1555.4 mbsf. On the way back to bottom, tight sections of hole had to be drilled through and reamed up and down (864875, 885887, 992, 1000, 10051024, 1093, 12851351, and 1365 mbsf). Torque of 400500 A and pump pressures up to 1700 psi were recorded while drilling out the upper two intervals. Increased torque and pump pressures also occurred from 1285 to 1351 mbsf. The typical drilling parameters while reaming back to the bottom of the hole were WOB = 512 klb, bit rotation = 80 rpm, torque = 175200 A, and circulation = 120 spm at 1200 psi. Sepiolite mud was circulated at 1300 mbsf (30 bbl) and 1514 mbsf (50 bbl).
Coring resumed at 1330 hr. Cores 81R to 84R were taken from 1555.4 to 1587.7 mbsf and we recovered 17.85 m (recovery = 55%). Coring parameters were WOB = 2035 klb, bit rotation = 6570 rpm, torque = 200300 A, and circulation = 120 spm at 20002200 psi. We circulated 30 bbl of mud every core. The AHC was used while cutting Core 82R. Recovery for this core was very low (0.94 m), due to a mechanical problem with the check ball that allowed the full pump rate to pass through the core barrel.
Because of the slow and apparently decreasing rate of penetration (ROP = ~3 m/hr for Cores 81R through 83R and 1.8 m/hr for Core 84R), we decided to retrieve the drill string so we could change the bit to one better suited to hard formations.
We deployed the wireline sinker bar to retrieve Core 84R at 0715 hr on 21 August. While lowering the sinker bar, the driller noticed that the top drive was leaking while he was circulating seawater (similar to the failure that occurred on 24 July). The situation changed drastically when the leak turned into a deluge. The driller immediately landed and secured the drill string on the rig floor elevators. Fortunately, the bit was off the bottom of the hole when this occurred, thus allowing the driller access to the tool joint on the 5.5-in drill pipe.
Once the drill pipe had been secured, our next concerns were to make sure the pipe didn't get stuck in the hole and to start retrieving the drill string as soon as possible. We needed to retrieve Core 84R, reestablish circulation, get the top drive out of the way, and then start pulling out of the hole.
Core 84R was pulled out of the pipe and recovered on the rig floor at 0830 hr on 21 August. The circulating head was then attached to the top of the drill pipe so that we could pump seawater to prevent the pipe from becoming stuck in the hole. The knobby drill pipe attached to the bottom of the top drive had to be removed before the top drive could be pulled out of the way. At 0830 hr, the 30-ft knobby was removed but the other 20-ft knobby had to be cut off because we could not reach the connection to loosen it. Once the knobbies were removed from the top drive, the top drive was rotated back away from the rig floor.
At 0930 hr, we started raising drill pipe out of the hole. The drill string encountered 50 klb of overpull and 25 klb of drag when raising and lowering the pipe through 1312 mbsf. Based on previous hole conditions when conducting wiper trips or pulling out of the hole, we were certain that circulation would be required to get out of the hole. Thus, we reinstalled the circulating head on the top of each stand of drill pipe. The pipe had to be worked up and down for ~2 hr before we could get through the interval from 1312 to 1323 mbsf. Tight hole conditions were encountered throughout the open hole and the pipe had to be worked up and down while circulating (pump rate = 30 spm at 300 psi). At 1930 hr, after 7.5 hr, the bit was raised back up inside the 16-in casing. We continued retrieving the drill string, which cleared the seafloor at 2055 hr and was back on the rig floor at 0455 hr on 22 August.
At 0500 hr, we began to disassemble the top drive to evaluate what had broken. A severe crack was found in the swivel-shaft box connection. The swivel shaft bears the entire weight of the 6147-m-long drill string (~680,000 lb). The crack extended ~60% around the circumference of the shaft. It is a major miracle that we did not lose the drill string, lose the hole, and severely damage the derrick. We replaced the swivel shaft and then reassembled and tested the top drive. The top drive was fully repaired at 1530 hr.
We attached a new bit (CC-4) to the BHA and lowered it to the seafloor. The subsea camera system was launched at 0000 hr on 23 August. Before we could reenter Hole 1276A, we performed routine preventative maintenance on the drill line (slip and cut; 00450215 hr) and we picked up 10 new stands of 5-in drill pipe, measured length, and checked interior diameter.
We began searching for Hole 1276A at 0245 hr, and we reentered it on the second attempt at 0355 hr. The end of the drill string still had a 25 m offset from the moonpool due to currents, and there was 3 m of heave.
By 0545 hr, the bit was near the base of the 16-in casing. The bit was lowered to 1006 mbsf without using the top drive to rotate or circulate; this was the first time it entered so freely. The top drive was installed when the bit encountered drag (2030 klb) at 1006 mbsf. The hole was redrilled from 1006 to 1587.7 mbsf (0930 hr on 22 August to 2045 hr on 23 August). Tight spots were encountered at 865, 990, 1006, 1070, and 1110 mbsf.
The RCB core barrel was dropped to start Core 85R at 2045 hr. While the core barrel was dropping to the bottom of the pipe, we continued to rotate and ream out the lowermost 8.5 m of hole.
Cores 85R and 86R were taken from 1587.7 to 1604.5 mbsf and recovered 14.59 m (recovery = 86%). We varied the coring parameters (WOB = 30 klb, bit rotation = 60 rpm, torque = 275 A, and circulation = 120 spm at 2300 psi) in an attempt to improve the rate of penetration (ROP = 2.32.7 m/hr), but there was no apparent improvement. Although the cores recovered alternating claystone and sandstone, there was no apparent change in drilling throughout the cored interval.
On 24 August, we cut Cores 85R to 90R from 6152.8 to 6200.8 mbrf (1592.81640.8 mbsf). Core recovery was 90%, and ROP was 3.0 m/hr. Coring parameters were WOB = 2530 klb, bit rotation = 60 rpm, torque = 225275 A, and circulation = 120 spm at 2300 psi. A sill was encountered at the bottom of Core 87R and in most of the upper part of Core 88R. Core 88R recovery was 97.1%, and ROP was 3.8 m/hr. The AHC was run on Cores 87R and 89R. Particulate tracers were deployed while cutting Core 89R. We continued to run full nonmagnetic core barrels on all odd-numbered cores.
On August 25, we cut Cores 90R to 95R from 6200.8 to 6242.9 mbsf (1640.81682.9 mbsf). Coring parameters were consistent: WOB = 25 klb, bit rotation = 60 rpm, torque = 225275 A, and circulation = 120 spm at 2300 psi. Core recovery was 86.8%, and ROP was 3.2 m/hr. The AHC was run while cutting Core 91R, but it then had to be shut down because of a hydraulic leak.
By midnight on 25 August, sea conditions had deteriorated and we encountered maximum heave of nearly 3 m. Throughout 26 August, we continued to experience heave of 1.52.0 m.
On 26 August, we cut Cores 95R to 100R from 6242.9 to 6290.8 mbsf (1682.91730.8 mbsf). Particulate tracers were deployed while cutting Core 99R. Coring parameters were WOB = 2030 klb, bit rotation = 6070 rpm, circulation = 120 spm at 2200 psi, and torque = 225350 A. Core recovery dropped from 80% in Core 96R to only 39% for Core 99R. For Cores 95R to 98R, ROP was 35 m/hr but then it dropped to 2.5 m/hr for Core 99R. We had to replace the lower 5-ft section of the nonmagnetic core barrel due to a cracked thread. This may have been damaged when we experienced the high heave late on 25 August or early 26 August. Although we held a constant WOB of 25 klb, WOB could have ranged from 12 to 37 klb because of the heave.
While we cut Core 100R, the circulating pressures dropped from 2200 to 17001800 psi and the torque increased to 400500 A. After advancing only 2 m in 4 hr, we decided to retrieve the core to see if we could tell why the penetration rate had dropped so dramatically. When we recovered Core 100R it was empty. We deployed another core barrel and it did not appear to latch in properly, so it was immediately retrieved. When it was back on the rig floor, we noticed that the core catcher fingers inside the core barrel were damaged, as if something had poked through the middle of the core barrel when it landed. At this point we believed that the bit must have been jammed with rocks, so we ran a bit deplugger down on the end of a core barrel. This did not appear to land properly, so we ran in again with another type of deplugger.
At this point we felt that whatever was jamming the bit might be dislodged, so we deployed the barrel for Core 101R at 1200 hr on 27 August. While cutting this core, we observed relatively smooth torque of ~300 A during the first hour. After this, the torque became very erratic (200500 A). After a 1-m advance (1731.11732.1 mbsf), we retrieved Core 101R at 1730 hr and it was empty.
It appeared that the bit was still jammed, so we decided to retrieve the drill string to replace the bit. We began raising the drill string at 1735 hr. The top drive was removed at 2245 hr when the bit was at 929 mbsf, but the circulating head had to be used from 865 to 813 mbsf to work through a section that had 3040 klb of drag.
The bit cleared the seafloor at 0155 hr on August 28 and was back on the rig floor at 1005 hr. Pieces of diabase (~1.5 m total length) were jammed into the flapper valve. From all of the operations described above (especially the damaged core catcher fingers), we inferred that this 1.5 m of diabase most likely fell out of the core barrel while we were retrieving Core 99R. In addition, the bit was severely damaged, with one cone missing and one cone completely seized up; the remaining two cones could be moved up and down nearly 1 in on the bit. The three remaining cones also had missing buttons and severe gouging, and the core guides were damaged.
Once the damaged bit and pieces of core were removed, we started to assemble a new C-7 bit to a boot basket sub (designed to catch broken pieces of metal in cavities along the outside of the sub) and then we lowered the pipe back to the seafloor. We changed the bit in order to minimize damage to the cutting structure by any remnants of the lost bit cone. We did not want to use conventional fishing tools or a magnet due to the poor hole conditions (it was taking 1020 hr to drill back to bottom of hole each time) and because we wanted to continue coring if possible. Before we reentered the hole, we performed preventative maintenance on the drill line (slip and cut).
We positioned the drill pipe at 4553 mbrf and started to maneuver the ship for reentering Hole 1276A at 2115 hr on 29 August. Once again we had significant difficulties reentering the hole due to currents (surface and deep), a completely sediment-covered reentry cone, and significant offset of the moonpool from the bit (~75 m). After 24 hr and 15 attempts, we reentered Hole 1276A at 2110 hr on 30 August. On this last and successful attempt, it at first appeared that we had missed the hole yet again because of difficulty in lowering the drill pipe. However, when we offset the ship back over the hole, the bit slipped in. It may be that we had actually stabbed the hole on previous attempts, but it seems that the offset had caused the bit to stick in the throat of the reentry cone.
When we lowered the bit into the hole, we encountered 30 klb of drag at 800 mbsf and had to raise the bit to 784 mbsf and install the top drive. It took ~28 hr to drill back down to the bottom of the hole at 1732.1 mbsf. We encountered tight spots at 800900, 1030, 1075, 1285, 1320, 1354, and 1410 mbsf. It took ~5 hr to ream out and clean up a tight spot at 1688 mbsf using a WOB of 20 klb and torque of 300400 A. Another difficult section of hole was encountered between 1720 and 1726.8 mbsf where we needed to use a WOB of 20 klb and torque of 300450 A.
We dropped the barrel for Core 102R at 0045 hr on 31 August and starting coring at 1732.1 mbsf. Because the cone that had broken off the previous bit was in the bottom of the hole, we used a low WOB and a slow rotation rate while lowering the bit 0.8 m to 1732.9 mbsf. It took 2 hr to advance these 2 m. Core 102R was recovered with only a single 2-cm-long piece of diabase. After cutting Core 103R from 1732.9 to 1734.9 mbsf (3 hr), it was retrieved and it was empty.
At 1030 hr on 31 August, we deployed a center bit to try to dislodge anything that might be jammed in the bit and preventing core recovery and to help break up any debris from the missing cone. At 1115 hr, when we were ready to lower the sinker bars to retrieve the center bit, torque increased to 400 A and pump pressure increased to 2750 psi. We did not deploy the sinker bars and we reamed the tight hole from 1726.8 to 1693.8 mbsf. While reaming we encountered drag of 40,000 lb, torque of 600 A, and pump pressures of 2750 psi.
At 1315 hr on 31 August, we retrieved the center bit, then lowered the bit back to the bottom of the hole (1693.81734.9 mbsf) and started cutting Core 104R. We took 4 hr to advance 2 m using WOB = 20 klb, bit rotation = 55 rpm, and circulation = 100 spm at 1750 psi. When Core 104R was recovered it was empty.
At 2215 hr, we dropped another center bit in a further attempt to dislodge anything that might be jammed in the bit and to help break up any debris from the missing cone. The center bit was recovered at 0115 hr on 1 September, and it showed significant damage, indicating that it had been grinding on the lost cone in the bottom of the hole. We deployed another center bit at 0115 hr; it was retrieved at 0400 hr, and it showed significant damage to the cutting structure, also indicating that it had been grinding on metal in the bottom of the hole.
At 0400 hr, we deployed yet another center bit. When the bit had been lowered back down to 1734.9 mbsf (2 m above the bottom of the hole), we lost the ability to rotate the drill string (even with torque of >550 A) and pump pressures climbed to 2600 psi. We started to try to work the pipe up and down with overpull of 130 klb and pump pressures up to 3000 psi. Rotation could not be reestablished even with torque of 600 A. By 0930 hr, we were able to remove only two singles of drill pipe and to raise the bit up to 1707.5 mbsf.
At 0930 hr, we lowered the wireline to remove the center bit. At 1045 hr, we removed the 30-ft knobby drill pipe below the top drive so that we would have more space in the rig to work the pipe up and down. We attempted to work the drill string from 1045 to 1230 hr with maximum overpulls of 160 klb, torque of 650 A, and pump pressures to 3000 psi, but we were unable to raise, lower, or rotate the drill string. At this time we decided to prepare for severing the drill string while we continued trying to free it.
At 1600 hr, the drill string began to move ever so slowly with WOB = 150 klb overpull and pump = 100 spm and 3000 psi. The hole problems seemed to be associated with a relatively unconsolidated sediment section recovered in Core 97R. This was the section of hole where we had problems on the previous day. By 1715 hr, we were able to remove another 9.6-m section of drill pipe. At this point, we could slowly raise the drill string with progressively less overpull. We were able to raise the drill string from 1698.1 to 1678.9 mbsf and remove two more singles of drill pipe with 80 klb overpull and pressures of 3000 psi.
We then raised the bit from 1678.9 to 1131.2 mbsf with WOB overpulls = 2025 klb, torque = 400500 A, and pump pressures = 2000 psi. We continued to raise the bit up into the base of the 16-in casing using the top drive. This required overpull of 1025 klb and torque of 400 A from 950 to 750 mbsf.
By this time, it was clear that Hole 1276A could not be deepened without casing the open hole above total depth, so we decided to terminate operations at the site. Before we pulled out of the hole, we wanted to identify the location of Hole 1276A with a marker buoy deployed from the subsea camera/sonar system. The bit cleared the seafloor at 0600 hr on 2 September. At 0615 hr we released the marker buoy 15 m from Hole 1276A at an azimuth of 225°. We then continued to retrieve the drill string and subsea camera/sonar. At 0910 hr we released the seafloor positioning beacon and it was back on deck at 1015 hr. Once the bit was back on the rig floor (1530 hr) we began the transit to proposed Site NNB-04A (Site 1277).
We began the transit from Site 1276 to Site 1277 at 1530 hr on 2 September. The beginning of the transit was delayed by 45 min because of a problem with the propulsion control system. Once the problem was resolved, the ship was under way at 1615 hr. After a transit of 21 nmi, we arrived at Site 1277 at 1800 hr. The ship was positioned over the site using GPS, and the thrusters were lowered by 1815 hr.
Our aim at Site 1277 was to sample the lowermost 30 m of sediment and then core into basement of a shallowly buried ridge before we had to leave for St. John's. Basement was estimated to be at 136 mbsf, so we decided to drill without coring to 100 mbsf and then core until time ran out.
We assembled a RCB BHA and started lowering it the seafloor at 1830 hr on 2 September. We attached the top drive and deployed a wash core barrel (0230 hr on 3 September) in preparation for drilling. The bit tagged the seafloor at 0300 hr at 4639.4 mbrf (Table T2). We drilled with the wash core barrel in place from 0 to 103.9 mbsf. Hard intervals were penetrated at 8589 and 97.5100 mbsf. The wash core barrel was retrieved, and it recovered 2.29 m of basement-related rocks, mostly basalts.
Cores 2R to 9R penetrated 103.9180.3 mbsf and recovered 29.34 m (recovery = 38%). Coring parameters were WOB = 20 klb, bit rotation = 70 rpm, torque = 175 A, and circulation = 70 spm at 800 psi. ROP was 5.5 m/hr. Cores 6R to 8R were taken without plastic core liners to try to minimize core jamming and, hence, low recovery that we were experiencing. The AHC was run while cutting Cores 6R and 8R. We circulated 70 bbl of mud after Core 6R. The final core of Site 1277, Leg 210, and ODP (Core 9R) was on the rig floor at 1200 hr on 4 September. The last core of ODP was truly spectaculara 10.5-m (recovery = 106%) core consisting of green and brown serpentinized peridotite with calcite veins.
We removed the top drive at 1300 hr and we began to retrieve the drill string. The seafloor positioning beacon was recovered and back on board at 1500 hr on 4 September. The bit cleared the rig floor at 2345 hr. The DCs were taken apart, and the ship was secured for transit at 0100 hr on 5 September.
We began the transit to St. John's at 0100 hr on 5 September. After traveling 374 nmi, the last scientific expedition of the Ocean Drilling Program concluded with the last line ashore at 1500 hr on 6 September 2003.
Next Section | Table of Contents